OpenStudy (anonymous):

Find the point on the graph of \[F(x)=1-x^{2}\],those are closest to (0, 0).

OpenStudy (anonymous):

the curve is following: http://www.wolframalpha.com/input/?i=y%3D+1+-+x%5E2

OpenStudy (anonymous):

I am not asking the curve lol.

hartnn (hartnn):

y^2+x^2 = r^2 minimize r. (1-x^2)^2 +x^2 = r^2 r^2 =1+x^4 + -x^2 diff, r and set it to 0 is this the approach ?

OpenStudy (anonymous):

Correct, keep going and type the answer.

hartnn (hartnn):

4x^2=2 x=+/- sqrt(1/2) y= 1/2 points (1/2,1/2),(-1/2,1/2)

OpenStudy (anonymous):

Great it is .

OpenStudy (anonymous):

Insert sqrt in the 1/2 and -1/2 of the X coordinates.

hartnn (hartnn):

oh, yes. typo.

OpenStudy (calculusfunctions):

You want the minimum (shortest) distance from the given point to the parabola. Let (x, y) be any point on the parabola. Using the distance formula, we obtain \[L =\sqrt{x ^{2}+y ^{2}}\] Now substitute y = 1 - x² to get \[L =\sqrt{x ^{2}+(1-x ^{2})^{2}}\] Squaring both sides yields \[L ^{2}=x ^{2}+(1-x ^{2})^{2}\] Differentiating both sides w.r.t. x yields \[2L \frac{ dL }{ dx }=2x +2(1-x ^{2})(-2x)\] For max or min, dL/dx = 0. Hence 0 = 2x - 4x(1 - x²) Factoring 2x from the right side yields 0 = 2x[1 - 2(1 - x²)] 0 = 2x(1 - 2 + 2x²) 0 = 2x(2x² - 1) \[x =0, x =\pm \frac{ \sqrt{2} }{ 2 }\] Thus \[F(\pm \frac{ \sqrt{2} }{ 2 })=\frac{ 1 }{ 2 }\] F(0) = 1 ∴ the points on the parabola closest to the origin are (-√2/2, 1/2), (0, 1) and (√2/2, 1/2).

OpenStudy (anonymous):

Great work, but (0, 1) is not the possible answer.

OpenStudy (calculusfunctions):

@Zekarias Yes you're right! I forgot to perform the first or second derivative test to determine which critical points are minimum. Sorry!

OpenStudy (anonymous):

Thanks for your time. I wish I gave you the medal.

OpenStudy (calculusfunctions):

That's OK! Thanks anyway.

hartnn (hartnn):

i don't mind if u give him......

OpenStudy (anonymous):

hey, you can give them all,,,, lol

OpenStudy (calculusfunctions):

Thanks guys! lol